OpenStudy (anonymous):

A dragster travels 1/4 mi in 6.7 s. Assuming that acceleration is constant and the dragster is initially at rest, what is its velocity when it crosses the finish line? Please show step by step...

OpenStudy (campbell_st):

so use \[average ..speed = \frac{distance..travelled}{time..taken}\]

OpenStudy (anonymous):

I tried but my answer was wrong.) I got 267 mi/h??

OpenStudy (campbell_st):

did you try feet/sec..?

OpenStudy (campbell_st):

and I think your answer is incorrect for the right formula

OpenStudy (campbell_st):

1/4 mile = 5280/4 = 1320 ft avg speed 1320/6.7 = ...

OpenStudy (campbell_st):

even using 0.25/6.7 will give miles/sec then multiply by 60 to get miles/min then multiply by 60 to get miles/hr

OpenStudy (anonymous):

the answer is given it's 269 mi/h ???

OpenStudy (campbell_st):

wow...